Vanessa is 1 year more than three times the age of Lori. The difference of their ages is 25. How old is Vanessa?
I know the answer, but I dont know how to work it out!! Thanks for your help.

Answer by nyc_function(2741)

Vanessa = 3L + 1
Let L = Lori's age
3L + 1 - L = 25
2L = 25 - 1
2L = 24
L = 24/2
L = 12
Lori is 12 years old.
Vanessa is 3L + 1 years old or 3(12) + 1 = 37 years old.